OpenStudy (anonymous):

3 square root of -125 =?

OpenStudy (anonymous):

do you mean the cubed root of -125?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

then -5 is the cubed root -5*-5*-5=-125

OpenStudy (anonymous):

\[\sqrt[3]{-125}?\]

OpenStudy (anonymous):

if yes than\[-1*\sqrt[3]{125}= i*5=5i\]